ΔH and ΔS for a certain reaction are 10 kJ mol^-1 and 10 J K^-1 mol^-1 respectively. What is ΔG in kJ mol^-1 at 400 K? 


1
Expert's answer
2022-03-17T14:56:27-0400

"\Delta{G}=\Delta{H}-T\Delta{S}=10\frac{kJ}{mol}-400\ K\times0.01\frac{kJ}{mol\cdot{K}}=6\frac{kJ}{mol}"


Answer: 6 kJ/mol